OpenStudy (anonymous):

Can someone help me graph this rational function? I'm having trouble understanding it. f(x) = two divided by quantity x squared minus two x minus three

OpenStudy (anonymous):

\[f(x) = 2 \div x ^{2} - 3\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

\[f(x)=\frac{2}{x^2-2x+3}\]

OpenStudy (anonymous):

woops sorry

OpenStudy (anonymous):

Yes satellite73 that's what it looks like

OpenStudy (anonymous):

your sure it is \(+3\) and not \(-3\) at the end right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

that's too bad, cause that would have been easier

OpenStudy (anonymous):

Since when do they like to make it easy for us? XD

OpenStudy (anonymous):

first thing to look for is the zeros of the denominator, because that will give you the vertical asymptotes but this denominator has not zeros

OpenStudy (anonymous):

true

OpenStudy (anonymous):

so the domain is all real numbers, no vertical asymptotes for this one that makes things harder

OpenStudy (anonymous):

but not too bad it should be clear that the horizontal asymptote is \(y=0\) right, or no?

OpenStudy (anonymous):

Does it have a range? Why is the horizontal asymptote y = 0?

OpenStudy (anonymous):

last question first the numerator is a fixed number, it is 2 the denominator will get larger and larger the bigger x gets so you will have a fraction with a numerator of 2, and a denominator that gets large without bound think of a fraction like \(\frac{2}{1,000,000}\)

OpenStudy (anonymous):

another way to check is that the denominator is a polynomial of degree 2 and the numerator is a polynomial of degree zero ( a constant) and since the degree of the top is smaller than the degree of the bottom, the vertical asymptote is \(y=0\) that is always true

OpenStudy (anonymous):

as for the range, if it is clear than \[x^2-2x+3\] has no zeros, it should also be clear that it is always positive and since \(2\) is positive that means \[\frac{2}{x^2-2x+3}\] is always positive so one part we know about the range is that \(y>0\)

OpenStudy (anonymous):

one more step, but let me know if i lost you yet

OpenStudy (anonymous):

No you are making sense to me. So the horizontal asymptote is y = 0, there are no vertical asymptotes here, the domain is all real numbers and the range is y >0. Is that all correct?

OpenStudy (anonymous):

all but the last part the domain is not \(y>0\) that is the lower bound for the domain, we have to find the upper one that is why i said "one more step" the domain will be \[0<y\leq something\]

OpenStudy (anonymous):

the numerator is fixed the denominator gets bigger as x goes to infinity or to minus infinity if we find the minimum value of the denominator that will give the maximum value of the function

OpenStudy (anonymous):

that is identical to finding the second coordinate of the vertex of \[y=x^2-2x+3\] first coordinate is \(-\frac{b}{2a}=-\frac{-2}{2}=1\) and the second coordinate is \[1-2+3=2\]

OpenStudy (anonymous):

so the very largest this can be is when the denominator is 2 and you would get \[f(1)=\frac{2}{2}=1\]

OpenStudy (anonymous):

so the domain is a measly \[0<y\leq 1\]

OpenStudy (anonymous):